OpenStudy (animelover91804):

Please help!! Will fan and medal!! Solve the following algebraic equation 6x - 10 = 3x + 8 You have to figure out what x is, but x has to be the same and so does the answer for both of the equations in order for this to be correct

OpenStudy (aihberkhan):

Okay first add 10 to both sides. When you do this you should get: \(6x = 3x + 18\)

OpenStudy (misssmartiez):

Simplifying 6x + -10 = 3x + 8 Reorder the terms: -10 + 6x = 3x + 8 Reorder the terms: -10 + 6x = 8 + 3x Solving -10 + 6x = 8 + 3x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3x' to each side of the equation. -10 + 6x + -3x = 8 + 3x + -3x Combine like terms: 6x + -3x = 3x -10 + 3x = 8 + 3x + -3x Combine like terms: 3x + -3x = 0 -10 + 3x = 8 + 0 -10 + 3x = 8 Add '10' to each side of the equation. -10 + 10 + 3x = 8 + 10 Combine like terms: -10 + 10 = 0 0 + 3x = 8 + 10 3x = 8 + 10 Combine like terms: 8 + 10 = 18 3x = 18 Divide each side by '3'. x = 6 Simplifying x = 6

OpenStudy (aihberkhan):

Then, subtract \(3x\) from both sides. When you do this you should get: \(3x = 18\)

OpenStudy (aihberkhan):

Now divide 3 from both sides. You should get: \(x = 6\)

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @Animelover91804

OpenStudy (misssmartiez):

If you need more help on how to solve x or how it has gotten that answer, click the link then help.

OpenStudy (anonymous):

@MissSmartiez , please don't copy and paste as it is against Open Study Policies and Terms

OpenStudy (aihberkhan):

Yes, Do not copy and paste. That is not your own work. Thank You. @MissSmartiez

OpenStudy (misssmartiez):

Though, sited where I have gotten the resource and I haven't called it my own.

OpenStudy (anonymous):

But you still copied and pasted

OpenStudy (misssmartiez):

I get where you are coming from.

OpenStudy (aihberkhan):

Yes, but you are not aloud to copy and paste because other people work hard creating their own answers. Dont do it next time. Thank You. @MissSmartiez

OpenStudy (anonymous):

Please look over the Terms and Agreements of Open Study Policies

OpenStudy (animelover91804):

I looked at that earlier

OpenStudy (aihberkhan):

I hope this helped @Animelover91804 I am always happy to help! :)

OpenStudy (misssmartiez):

Did it help you Anime lover, when you looked at it earlier?

OpenStudy (animelover91804):

Thank you guys for the help!

OpenStudy (aihberkhan):

No problem! Anytime!! :) @Animelover91804